Assume that assets and liabilities increased by Br.240,000, and Br. 120,000 respectively during a given year.

Assume the following additional particulars further

▪️ Revenues generated during the year…. Br.80,000

▪️ Additional investment made by the owner during the year ………. Br. 70,000

▪️ Amount withdrawn by the owner during the year…. $10,000

Required: correct if the next calculation of the expense based on the above question.

if it is correct show every step of the calculation and write the explanation.

revenue..................8,000

Expense..................267050=8000+240000-120000+E

267050=236000+E

EXPENSE= 31050


Increase in Asset Gain .................240000

increase in Liability...................... (120000)

Net income...................................267050


1
Expert's answer
2022-01-25T17:42:58-0500

From logic, change in Assets- change in liabilities= Net Income

Assets = Liabilities + Equity

Equity= capital + Net income - withdrawal

Net Income= 240000- 120000- 10000= 110,000

Net income= Revenue- Expenses

110,000= (80,000+70,000)- Expenses

Expenses= 150,000- 110,000

= 40,000